OpenStudy (anonymous):

PLEASE HELP ME (kind of in a hurry) Choose the correct answer. Allan and Dave bowl together and their combined total score for one game was 375 points. Allan’s score was 60 less than twice Dave’s. What were their scores? Which is a system of equations to model the problem if x represents Dave’s score and y represents Allan’s score? x + y = 60 y = 2x – 375 x + y = 375 y = 2x – 60 x + y = 375 y = 2x + 60 x – y = 375 y = 2x – 60

OpenStudy (texaschic101):

x + y = 375 y = 2x - 60

OpenStudy (quin100):

same here

OpenStudy (anonymous):

thanks

OpenStudy (quin100):

yup

OpenStudy (anonymous):

i agree

OpenStudy (texaschic101):

great minds think alike :)

OpenStudy (ageta):

answers could be x + y = 60 y = 2x – 375 x + y = 375 y = 2x – 60 x + y = 375 y = 2x + 60 x – y = 375 y = 2x – 60

OpenStudy (texaschic101):

x + y = 375 y = 2x - 60 x + 2x - 60 = 375 3x - 60 = 375 3x = 375 + 60 3x = 435 x = 435/3 x = 145 x + y = 375 145 + y = 375 y = 375 - 145 y = 230 Dave (x) scored 145 Allen (y) scored 230